# Observational Signatures of Simpson-Visser Wormholes

### v1

@@ IV Hot Spot
Figure 9: Snapshots of different-side hot spots for wormholes with a single (Left) and double (Right) photon spheres, captured at the highest (Top), second-highest (Middle) and third-highest (Bottom) magnitude peaks.
The light curve and centroid motion of the hot spot located on the other side are presented in FIG. 8. Similar to the same-side case, the single-photon sphere wormhole exhibits a light curve with a prominent peak and a secondary peak. Consistent with this, the snapshots in the left panel of FIG. 9 reveals that the highest and second-highest peaks are dominated by the $n=1$ and $n=2$ images, respectively. Furthermore, due to image confinement within the critical curve, the centroid motion is restricted to a smaller range.
Intriguingly, the hot spot orbiting the double-photon sphere wormhole displays three distinct peaks of comparable heights in the light curve, as shown in the upper-right panel of FIG. 8. The corresponding snapshots in the right panels of FIG. 9 reveal that the highest peak is dominated by a combination of $n=2$ and $3$ images, the second-highest peak by $n=1$ and $2$ images, and the third-highest peak by $n=3$ and $4$ images. Consequently, the presence of higher-order images leads to a more irregular centroid motion in this case.
@@ V Conclusions
This paper investigates the images of traversable wormholes illuminated by celestial spheres and hot spots residing in two separate universes. When the celestial sphere and hot spot reside on the same side as the observer, their images closely resemble those of a black hole. However, significant differences arise when the light sources are positioned on the different side relative to the observer:
@@
  • Different-side scenario: The celestial sphere always forms images within the critical curve, with compressed images asymptotically approaching the critical curve. The hot spot manifests as two distinct image tracks for the single-photon sphere case, while the double-photon sphere case exhibits four distinct image tracks. In the single-photon sphere case, the two tracks again lead to two peaks in the light curve. Conversely, the combination of the four tracks in the double-photon case results in three peaks in the light curve.
By analyzing these image characteristics, we can gain valuable insights into the optical signatures of light sources near wormholes. This understanding has the potential to not only differentiate between wormholes and black holes but also discriminate between single- and double-photon sphere wormholes. While observing images from the other side of the wormhole presents an intriguing prospect, it also poses greater challenges. Therefore, higher resolution instruments, such as the next-generation Very Long Baseline Interferometry, are also keenly expected.
@@ Acknowledgements.
We are grateful to Yiqian Chen for useful discussions and valuable comments. This work is supported in part by NSFC (Grant No. 12105191, 12275183, 12275184 and 11875196).

### v3

@@ IV Hot Spot
Figure 9: Snapshots of the hot spot orbiting wormholes on the side opposite to the observer, for wormholes with a single (Left Column) and double (Right Column) photon spheres. The snapshots are captured at the highest (Top Row), second-highest (Middle Row) and third-highest (Bottom Row) magnitude peaks.
The corresponding light curve and centroid motion of the hot spot are presented in the top and bottom rows of FIG. 8, respectively. Similar to the same-side case, the single-photon sphere wormhole exhibits a light curve with a prominent peak and a secondary peak. Consistent with this, the snapshots in the left panel of FIG. 9 reveals that the highest and second-highest peaks are dominated by the $n=1$ and $n=2$ images, respectively. Furthermore, due to the confinement of the images within the critical curve, the centroid motion is restricted to a smaller range. Intriguingly, the hot spot orbiting the double-photon sphere wormhole displays three distinct peaks of comparable heights in the light curve, as shown in the upper-right panel of FIG. 8. The corresponding snapshots in the right panels of FIG. 9 reveal that the highest peak is dominated by a combination of $n=2$ and $3$ images, the second-highest peak by $n=1$ and $2$ images, and the third-highest peak by $n=3$ and $4$ images. Consequently, the presence of more higher-order images leads to a more irregular centroid motion in this case.
@@ V Conclusions
This paper investigates the images of traversable wormholes illuminated by celestial spheres and hot spots. Two spacetimes are connected at the wormhole throat, which allows light rays to travel from one side to the other one. When the celestial sphere and hot spot reside on the same side as the observer, their images closely resemble those of a black hole. However, significant differences arise when the light sources are positioned on the different side from the observer. Specifically, we found that
@@
  • Different-side scenario: The celestial sphere always forms images within the critical curve, with compressed higher-order images asymptotically approaching the critical curve. The hot spot manifests as two distinct image tracks for the single-photon sphere case, while the double-photon sphere case exhibits four distinct image tracks. In the single-photon sphere case, the two tracks again lead to two peaks in the light curve. Conversely, the four tracks in the double-photon case result in three peaks in the light curve.
By analyzing these image characteristics, we can gain valuable insights into the optical signatures of light sources near wormholes. This understanding has the potential to not only differentiate between wormholes and black holes but also discriminate between single-photon sphere and double-photon sphere wormholes. While observing images from the other side of the wormhole presents an intriguing prospect, it also poses greater challenges. Therefore, higher resolution instruments, such as the next-generation Very Long Baseline Interferometry, are also keenly expected.
@@ Acknowledgements.
We are grateful to Tianshu Wu for useful discussions and valuable comments. This work is supported in part by NSFC (Grant No. 12105191, 12275183, 12275184 and 11875196).